# On the Identifiability of Steering Vectors in Large Language Models

### v1

@@ Abstract
Activation steering methods, such as persona vectors, are widely used to control large language model behavior and increasingly interpreted as revealing meaningful internal representations. This interpretation implicitly assumes steering directions are identifiable and uniquely recoverable from input–output behavior. We formalize steering as an intervention on internal representations and prove that, under realistic modeling and data conditions, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we validate this across multiple models and semantic traits, showing orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes. However, identifiability is recoverable under structural assumptions including statistical independence, sparsity constraints, multi-environment validation or cross-layer consistency. These findings reveal fundamental interpretability limits and clarify structural assumptions required for reliable safety-critical control.
On the Identifiability of Steering Vectors in Large Language Models
Sohan Venkatesh  and Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru {sohan1, ashish}.mitblr2022@learner.manipal.edu
@@ 1 Introduction
Persona vector steering has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations. Empirically, such vectors can shift model outputs along interpretable dimensions such as politeness, political ideology or truthfulness, suggesting that representational alignment might afford fine-grained behavioral control without retraining ((Zou et al., 2023; Rimsky et al., 2024; Turner et al., 2023)). This line of work is closely connected to broader efforts in representation engineering and activation editing, where linear directions in activation space are used to modulate model behavior along semantic axes ((Elhage et al., 2022; Turner et al., 2024)).
@@
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors—for example, "the politeness direction" or "the honesty direction"—and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints ((Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019)). Nonlinear ICA and causal representation learning further emphasize that recovering latent factors from high-dimensional observations is generically impossible without auxiliary information or strong inductive biases ((Hyvarinen and Morioka, 2017; Khemakhem et al., 2020; Ahuja et al., 2022)).
This raises a fundamental question for alignment and interpretability research: when does representational alignment genuinely afford reliable behavioral control and when does it merely exploit underdetermined or spurious correlations? Existing evidence from probing and representation analysis suggests that even seemingly meaningful linear directions can reflect artifacts of the measurement procedure rather than uniquely defined semantic variables ((Hewitt and Liang, 2019; Ravfogel et al., 2020; Elazar et al., 2021; Belinkov, 2022)). In the context of activation steering, these concerns become particularly acute, because the resulting vectors are often used not only for analysis but also for safety-relevant control interventions. Understanding identifiability is therefore critical for several reasons:
Alignment affordances.
If persona vectors are not identifiable, then representational alignment may provide only heuristic control rather than principled intervention. Characterizing when steering vectors are unique clarifies when alignment interventions can be trusted and when they should instead be viewed as exploiting one of many behaviorally equivalent directions.
Interpretability validity.
When multiple incompatible vectors produce identical observable behavior, claims that a specific vector "represents" a semantic concept become scientifically underdetermined. Identifiability theory distinguishes well-grounded interpretability claims from artifacts of measurement and projection ((Elazar et al., 2021; Marks and Tegmark, 2023)).
Robustness and safety.
Non-identifiable steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting. For safety-critical applications—where steering vectors may be used to enforce norms or prevent harmful behavior—understanding identifiability limits is essential to avoid brittle or misleading forms of control.
Methodological design.
Identifiability theory clarifies which experimental protocols provide meaningful evidence and which require additional structure to support reliable conclusions. In particular, it highlights when interventions on internal activations can be interpreted causally and when they merely reparameterize an equivalence class of representations ((Schölkopf et al., 2021; Ahuja et al., 2022)).
@@
In this work, we provide, to our knowledge, the first formal identifiability analysis of persona vector steering. Our contributions are threefold. First, we prove that under standard observational regimes, persona vectors are generically non-identifiable due to null-space ambiguity in the model’s input-output Jacobian (Proposition 1). Specifically, under white-box single-layer access, infinitely many geometrically distinct steering directions induce identical observable behavior. Second, we characterize sufficient structural conditions under which identifiability can be recovered including statistical independence constraints, sparsity priors, multi-environment access and cross-layer consistency (Proposition 2). Third, we demonstrate empirically that contemporary steering operates in the non-identifiable regime across multiple models and semantic traits, with vectors perturbed by orthogonal components achieving near-equivalent efficacy. Crucially, our results are not purely negative. By explicitly characterizing both the limits and the affordances of persona vector steering, we provide principled guidance for when representational alignment can be trusted and which structural assumptions enable reliable control in safety-critical applications.
@@ 2 Related Work
Our work formalizes persona steering as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
Causal and latent variable identifiability.
Classical results show that latent variable models are generically non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations. Nonlinear ICA methods (Khemakhem et al., 2020; Hyvarinen and Morioka, 2017) provide theoretical foundations for recovering latent variables using temporal or auxiliary information, which standard steering methods do not exploit.
Probing and representation learning.
Pimentel et al. (2020) and Ravfogel et al. (2020) demonstrate that probing classifiers can succeed for trivial reasons unrelated to the presence of target information. Elazar et al. (2021) show that removing information via projection is ill-defined without identifiability guarantees. Belinkov (2022) provides a comprehensive survey of probing methods and their limitations. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022; Marks and Tegmark, 2023) suggests that concepts correspond to directions in activation space but lacks formal identifiability guarantees.
Activation editing in LLMs.
Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024), activation patching (Meng et al., 2022) and causal tracing (Wang et al., 2022) manipulate model internals to control behavior. While empirically successful, these works generally do not address whether the resulting control directions are uniquely determined. Related work on linear mode connectivity (Entezari et al., 2021) and neural network reparameterization (Dinh et al., 2017) reveals symmetries in neural representations that motivate our null-space analysis.
Steering and control vectors.
Early work on steering vectors (Li et al., 2022; Turner et al., 2023; Tigges et al., 2023) develops techniques for extracting control directions from contrastive prompt pairs. Burns et al. (2022) investigate discovering latent knowledge in language models. These methods assume that extracted vectors correspond to uniquely determined semantic factors—an assumption we formally examine.
Mechanistic interpretability.
Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
Compressed sensing and sparse recovery.
The theoretical framework of compressed sensing (Candès and Wakin, 2008; Donoho, 2006) provides conditions under which sparse vectors can be uniquely recovered from linear measurements. These results directly inform the sparsity-based identifiability conditions in Proposition 2.
@@ 3.2 Observational Regimes
$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})].$
@@ 3.3 Linear Approximation and Nonlinear Case
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v,$
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|{h\ell(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian and $V$ is the vocabulary size.
@@ 5 Main Results
We now state our main theoretical results, characterizing when observational conditions afford identifiable persona vectors and when they do not.
5.1 Proposition 1: Non-Identifiability Without Structural Constraints
@@
We establish non-identifiability via two complementary arguments: reparameterization symmetry and null-space ambiguity.
Reparameterization symmetry. Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and define the reparameterized representation $h^{\prime}\ell(x)=T(h\ell(x))$. If the subsequent layers can be equivalently expressed as $F_{\ell\to L}(h_{\ell})=F^{\prime}{\ell\to L}(T(h\ell))$, then using $(h_{\ell},v)$ or $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=T(v)$ is observationally indistinguishable.
For overparameterized networks, many such reparameterizations exist that leave the input-output function invariant. Consider $T(h)=Ah+b$ where $A\in\mathbb{R}^{d\times d}$ is invertible. This transformation can be absorbed into the parameters of layer $\ell+1$ through weight reparameterization. While we cannot explicitly construct these transformations for a given frozen model without retraining, their existence follows from the network’s inherent symmetry structure. Let $A$ be any invertible matrix with $A\neq cI$ for any scalar $c$. Then $v^{\prime}=Av$ is not proportional to $v$, yet produces equivalent observations after appropriate reparameterization. Since there are infinitely many such $A$, there exist infinitely many equivalent $v^{\prime}$.
Null-space ambiguity (primary mechanism). Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$ and thus produces identical linearized outputs. This establishes non-identifiability even without invoking global reparameterization symmetries.
Corollary 1.1.
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$.
@@
The null space $\ker(J_{\ell})$ is typically high-dimensional in practice. For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$. In modern language models, $V\approx 50000$ and $d\approx 4000$, so max rank is $d$. However, the output distribution lies on a low-dimensional manifold, causing $\text{rank}(J_{\ell})\ll d$ in practice, consistent with observations that neural network representations concentrate on low-dimensional subspaces (Maennel et al., 2018; Li et al., 2018). This yields $\dim(\ker(J_{\ell}))=d-\text{rank}(J_{\ell})\gg 0$. This result establishes generic non-identifiability under local linear approximation.
5.2 Proposition 2: Identifiable Regimes Under Structural Assumptions
Proposition 2.
Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
  • Statistical Independence (ICA). If the latent persona $z=(z_{1},\ldots,z_{k})$ has independent components and $h_{\ell}=Az+\epsilon$ where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix, then under non-Gaussianity of $z_{i}$ and sufficient observations, AA (and hence the columns $v_{i}$) can be recovered up to permutation and scaling (Comon, 1994; Hyvärinen and Oja, 2000).
  • Sparsity Constraints. If the true persona vector $v$ is sparse (i.e., $|v|_{0}\leq s\ll d$) and we observe the effect of steering on multiple outputs, then under restricted isometry properties, $v$ can be uniquely recovered via $\ell_{1}$ minimization(Candes and Tao, 2005).
  • Multi-Environment or Interventional data. If we observe the same persona $z$ across multiple contexts (prompts, models or layers) where the spurious correlations change but the true signal remains stable, then techniques from causal representation learning allow identification of invariant factors. (Peters et al., 2017; Ahuja et al., 2022).
  • Cross-layer Consistency. If we assume persona vectors have consistent geometric relationships across multiple layers (e.g., $v_{\ell}\approx W_{\ell}v_{\ell-1}$ for known or estimable $W_{\ell}$), then the overdetermined system provides additional constraints that break symmetries.
Interpretation.
Proposition 2 serves as a natural theoretical extension of Proposition 1: by characterizing sufficient conditions under which identifiability could be recovered–such as statistical independence (ICA), sparsity priors, multi-environment access or cross-layer consistency. Proposition 2 clarifies precisely which structural assumptions are necessary to break the symmetries that cause non-identifiability.
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5.1) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
@@ 6.1 Experimental Setup
We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales:
  • Qwen2.5-3B-Instruct: 24 layers, hidden dimension $d=2048$, layer $\ell=12$ (mid-network)
  • Llama-3.1-8B-Instruct: 32 layers, hidden dimension $d=4096$, layer $\ell=16$ (mid-network)
For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., ) in steering research. Middle layers balance semantic abstraction with steerability: early layers encode low-level features (tokens, syntax), while late layers specialize for next-token prediction.
@@
We test three semantic traits spanning distinct dimensions:
  • Formality: Formal versus informal style
  • Politeness: Polite versus rude social register
  • Humor: Humorous versus serious content
This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
@@
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. This provides a conservative test of observational equivalence: if vectors are distinguishable semantically, they are certainly non-identical in the full distributional sense.
For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.

### v2

# On the Non-Identifiability of Steering Vectors in Large Language Models

@@ Abstract
Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations. This interpretation assumes steering directions are identifiable and uniquely recoverable from input–output behavior. We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we show that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits. Critically, we show that the non-identifiability is a robust geometric property that persists across diverse prompt distributions. These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions. [^1] Code: https://github.com/sohv/non-identifiability
On the Non-Identifiability of Steering Vectors in Large Language Models
Sohan Venkateshthanks: Correspondence: soh.venkatesh@gmail.com     Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru
@@ 1 Introduction
Persona vector steering (Chen et al., 2025) has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations. This line of work is closely connected to broader efforts in representation engineering (Zou et al., 2023) and activation editing (Rimsky et al., 2024; Turner et al., 2023), where linear directions in activation space are used to modulate model behavior along semantic axes (Elhage et al., 2022; Turner et al., 2024).
@@
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
This raises a fundamental question: when does representational alignment afford reliable behavioral control? Understanding identifiability is critical for several reasons. First, it determines when persona vectors provide principled causal interventions versus merely heuristic control that exploits one of many behaviorally equivalent directions. Second, it clarifies when claims that a specific vector “represents” a semantic concept are scientifically well-grounded rather than artifacts of measurement and projection (Elazar et al., 2021; Marks and Tegmark, 2023). Third, it reveals when steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting. Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
Figure 1: Geometric illustration of steering vector non-identifiability. Two geometrically distinct vectors, the original steering vector $v$ (left) and a perturbed version $v+v_{\perp}$ where $v_{\perp}\in\mathrm{ker}(J)$ (right), produce identical observable outputs. The null-space component (red, unobservable region) can be arbitrarily modified without changing outputs, demonstrating the existence of large equivalence classes of behaviorally indistinguishable steering directions.
@@
We make three contributions. First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity. Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits. Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than an artifact of limited prompt diversity. These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
@@ 2 Related Work
Our work formalizes persona steering  [^2] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025). as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
Causal and latent variable identifiability. Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
Probing and representation learning. Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
Activation editing in LLMs. Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined. Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
Robustness and invariant representations. Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations. Methods for learning invariant representations (Arjovsky et al., 2019; Krueger et al., 2021) aim to extract features stable across domains. Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
Mechanistic interpretability. Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
@@ 3.2 Observational Regimes
$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})]$ (3)
@@ 3.3 Linear Approximation and Nonlinear Case
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (4)
where $J_{\ell}(x)=\left.\frac{\partial o}{\partial h_{\ell}}\right|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ denotes the Jacobian of the output and $V$ is the vocabulary size.
@@ 5 Main Results
We now state our main theoretical result, proving that persona vectors are fundamentally non-identifiable under standard observational regimes.
@@
We establish non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
Null-space ambiguity (primary mechanism). Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$, and thus produces identical linearized outputs. This establishes non-identifiability even without invoking global reparameterization symmetries.
Reparameterization symmetry. Neural networks exhibit inherent parameter symmetries arising from overparameterization. Consider an invertible linear transformation $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$. For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, the computation $h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ can be equivalently expressed through reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}Ah_{\ell}+b_{\ell+1}\right)$ (7)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}$
Under this reparameterization, steering with $v^{\prime}=Av$ on the transformed representation $h^{\prime}_{\ell}=Ah_{\ell}$ produces identical outputs to steering with $v$ on $h_{\ell}$. For any invertible matrix $A\neq cI$ (not a scalar multiple of identity), we obtain $v^{\prime}\not\propto v$, yet these vectors are observationally equivalent. Since the space of such matrices has dimension $d^{2}-1$, infinitely many distinct reparameterizations exist.
Practical vs. theoretical relevance: The reparameterization argument establishes non-identifiability as a fundamental property of the model class, independent of any specific instantiation. While these symmetries cannot be exploited in frozen models without modifying weights, their existence proves that identifiability fails even in principle. For analyzing deployed models, the null-space mechanism provides direct, operationally relevant non-identifiability.
Corollary 1.1 (Null-space equivalence).
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$, since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
@@
The null space $\ker(J_{\ell})$ is typically high-dimensional in practice. For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$. In modern language models, $V\approx 50{,}000$ and $d\approx 4{,}000$, so the maximal rank is $d$. However, empirical evidence shows that (1) neural network representations exhibit low intrinsic dimensionality (Li et al., 2018; Maennel et al., 2018), and (2) output sensitivity concentrates in a small number of Hessian directions (Ghorbani et al., 2019). These findings suggest the Jacobian $J_{\ell}$ has effective rank substantially smaller than $d$, implying a high-dimensional null space:
$\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$ (8)
@@ 6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
@@ 6.1 Experimental Setup
We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$). For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
@@
We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor. This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
@@
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.